OpenStudy (anonymous):

Please help! GI bisects angle DGH so that angle DGI is x – 3 and angle IGH is 2x – 13. Find the value of x

OpenStudy (anonymous):

Angles DGI and IGH will be equal since GI is the bisector therefore, x-3 = 2x - 13 10 = x 10-3 = 7 2(10) - 13 = 7. 7 = 7. You can check my answer it should be right. (I checked it already) Weird the website just went down...

OpenStudy (anonymous):

Thank you so much! And they were fixing a bug.

OpenStudy (anonymous):

ok